OpenStudy (anonymous):

Prove that (1 + secx)(1 - cosx) = sinxtanx??

OpenStudy (amistre64):

write everything in terms of sin and cos to see things better

OpenStudy (anonymous):

so 1 - cosx + secx - cosxsecx = secx - cosx. Now what??

OpenStudy (anonymous):

sec=cos/sin

OpenStudy (anonymous):

no, sec is 1/cos

OpenStudy (anonymous):

hhaha

OpenStudy (amistre64):

i was thinking more of: \[1-cos+sec-cos.sec=sin.tan\] \[1-cos+sec-cos/cos=sin.sin/cos\] \[1-cos+sec-1=sin^2/cos\] \[-cos+1/cos=sin^2/cos\] for starters

OpenStudy (anonymous):

but we don't have to do that. we have to work from one side, and expand that to get the other side. It has to be proven.

OpenStudy (amistre64):

this is working one side; I havent changed the left at all except to see it better

OpenStudy (amistre64):

and by left i mean right lol

OpenStudy (amistre64):

\[-cos+1/cos=sin^2/cos\] \[(-cos^2+1)/cos=sin^2/cos\] \[(1-cos^2)/cos=sin^2/cos\] \[sin^2/cos=sin^2/cos\]

OpenStudy (anonymous):

oh silly me it is a product!! i thought it was a fraction...

OpenStudy (anonymous):

but we are not supposed to even supposed to do ANYTHING with the other side. It's like we're not supposed to know the other side.

OpenStudy (anonymous):

^that's right :P

OpenStudy (amistre64):

your requirements are absurd

OpenStudy (amistre64):

your not spose to know tha tan = sin/cos? your not spose to know that sin.sin = sin^2? your not spose to know what; any trig relations and identities? but somehow transform them into each other?

OpenStudy (anonymous):

not that, yaar. Imagine they give you a question, saying expand ....., leaving it terms of sin/cos/tan. SO we just do that. we expand. but how??

OpenStudy (anonymous):

\[(1+\frac{1}{a})(1-a)=\frac{1}{a}-a=\frac{1-a^2}{a}\] now replace a by cosine and you should get it

OpenStudy (amistre64):

since the right side isnt changed; just revert it back into its origanl form and and leave it there; but the process is still the same ...

OpenStudy (anonymous):

it is much easier to do the algebra without the sines and cosines. then put them in at the end.

OpenStudy (anonymous):

but how does that give us sinxtanx??

OpenStudy (anonymous):

but what amistre said is important. it in not a matter of leaving one side untouched. if you can change one side, you can change it back. if you have to hand it in in a certain form, fine, but if you can alter one side you can presumably go backwards

OpenStudy (amistre64):

what? how does sin.sin/cos = sin tan?

OpenStudy (amistre64):

by definition of what sin and tan ARE lol

OpenStudy (anonymous):

ok we did the algebra part, now comes the trig part. replace a by cosine and get \[\frac{1-\cos^2(x)}{\cos(x)}\] now come the only trig step, writing \[1-\cos^2(x)=\sin^2(x)\] that is the one step that is trig

OpenStudy (amistre64):

can know that 1-cos^2 = sin^2 sat; its forbidden knowledge ;)

OpenStudy (anonymous):

I GOT IT. I love you guys :D

OpenStudy (anonymous):

so you get \[\frac{1-\cos^2(x)}{\cos(x)}=\frac{\sin^2(x)}{\cos(x)}=\frac{\sin(x)}{\cos(x)}\times \sin(x)\]

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

and we love you!

OpenStudy (anonymous):

okayy, another question. given that tanx = p, find an expression, in terms of p, for cosec^2 x. GO.

OpenStudy (amistre64):

gone!! ;)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh i see what they want

OpenStudy (anonymous):

so is it 1 + 1/p^2 ??

OpenStudy (anonymous):

i don't think so let me draw a triangle. you should draw one too

OpenStudy (anonymous):



OpenStudy (anonymous):

there is my triangle, the triangle where tan(x) = p

OpenStudy (anonymous):

now you need the hypotenuse which you find by pythagoras

OpenStudy (anonymous):